OpenStudy (anonymous):

Louann and Carla received equal scores on a test made up of multiple choice questions and an essay. Louann got 14 multiple choice questions correct and received 24 points for her essay. Carla got 16 multiple choice questions correct and received 14 points for her essay. How many points was each multiple choice question worth?

OpenStudy (jojokiw3):

Okay. Nice profile pic. 16 - 14 = 2. 24 - 14 = 10. Now since they both got the same store, 2 multiple choice questions must have the value of 10 points. 10/2 = 5. Each multiple choice question is worth 5 points.

OpenStudy (anonymous):

thanks bro nice profile 2